OpenStudy (anonymous):

What is the indefinite integral of (x^5-3)^2? Thanks in advance!

OpenStudy (anonymous):

first expand the expression

myininaya (myininaya):

\[\int\limits_{}^{}(x^{10}-6x^5+9) dx\]

myininaya (myininaya):

i think this form you should know how to do just use \[\int\limits_{}^{}x^n dx=\frac{x^{n+1}}{n+1}+c\]

myininaya (myininaya):

you have three parts just decide what n is in use the formula for each of those three parts

OpenStudy (anonymous):

\[x ^{11}/11-6x ^{7}/7+9x\]

myininaya (myininaya):

almost luckey

OpenStudy (anonymous):

isn't it 6x^6/6?

myininaya (myininaya):

very good! :) don't forget the minus sign in front of and you are awesome! :)

OpenStudy (anonymous):

yeah plus a constant

OpenStudy (anonymous):

One more if you guys don't mind? This one is a little more complicated... indefinite integral of xsqrt(2x+9)?

OpenStudy (anonymous):

with steps if you can.

myininaya (myininaya):

let u=2x+9 => du=2 dx => du/2 =dx (u-9)/2=x \[\int\limits_{}^{}\frac{u-9}{2} \sqrt{u} \frac{du}{2}\]

myininaya (myininaya):

that should help \[\frac{1}{4}\int\limits_{}^{}(u \cdot u^\frac{1}{2}-9u^\frac{1}{2}) du\]

OpenStudy (anonymous):

Thanks!